OpenStudy (anonymous):

2a^3+8a^2+11a-8/a+2= 2a^2+4a+3-14/a+2 ?

OpenStudy (lgbasallote):

cross multiply a + 2...it will cancel right? rewrite the equation now withoout a+ 2..you might have some ideas

OpenStudy (anonymous):

i was asking if 2a^2+4a+3-14/a+2 was the answer

OpenStudy (lgbasallote):

ohh..be clearer man =))) so howd you get 4a + 3 - 14?

OpenStudy (anonymous):

a+2 / 2a^3+8a^2+11a-8 a*2a^2=2a^3 and so on with the rest

OpenStudy (anonymous):

\[2 a^3+6 a^2+7 a+\frac{6}{a}-3=0 \]